OpenStudy (anonymous):

Solve: 7(x − 1) = 5(x + 2) + 1 9 18 2 4

OpenStudy (anonymous):

7(x - 1) = 5(x + 2) + 1 => 7x - 7 = 5x + 10 + 1 => 2x - 7 = 11 => 2x = 18 => x = 9